# Fast-Revisit Coverage Path Planning for Autonomous Mobile Patrol Robots Using Long-Range Sensor Information

### v1

@@ Abstract
In this paper, we propose the novel Fast-Revisit Coverage Path Planning (FaRe-CPP) algorithm using a greedy heuristic approach to propose waypoints for maximum coverage area and a random search-based path optimization technique to obtain a path along the proposed waypoints with minimum revisit time. We evaluated the algorithm in a simulated environment using Gazebo and a camera-equipped TurtleBot3 against a number of existing algorithms. Compared to their average revisit times and path lengths, our FaRe-CPP algorithm approximately showed a 45% and 40% reduction, respectively, in these highly relevant performance indicators.
@@
Ensuring minimum revisit times of any given waypoint is crucial for certain application, for instance, safety patrol on hazardous substance sites like garbage dumps and recycling centers. Typically, in such sites, spontaneous fires are an imminent hazard because of inflammable substances or self-ignition because of the heat produced during composting. Therefore, fire incidents happen frequently [3, 4] and there is great interest by waste disposal enterprises to prevents such indicents. A safety patrol robot that is able to detect fires autonomously, and potentially even applies extinguishing agents directly, calls the fire department and keeps the fire at bay until their arrival. For such kind of patrols, minimum revisit times with maximum coverage are essential to prevent fires from getting out of control.
In this paper, we present a novel Coverage Path Planning algorithm with focus on fast revisit times for a single mobile patrol robot. We name the algorithm Fast-Revisit Coverage Path Planning (FaRe-CPP). Utilizing a pre-recorded 2D occupancy grid map, our approach updates the surveilled region within the field of view of the relevant onboard sensors at each step, subsequently creating waypoints that maximize the coverage area. Initially generated with a greedy heuristic for maximum perception potential, these waypoints are optimized using the GRASP method [5] to minimize path length while preserving coverage information. The combination of waypoint generation subsequent optimization ensures maximum overall area coverage, with no path overlap and minimal path length, resulting in minimal revisit times.
A Python implementation of FaRe-CPP with ROS111Robot Operating System, https://ros.org connectivity is publicly available on https://github.com/hcr-lab/fare-cpp
@@ II Related Work
II-A Early foundations for Coverage Path Planning
@@
Coverage Path Planning is essential for robot applications like vacuum cleaning [6], radiation monitoring [7], bridge inspection [8], emergency evacuation [3], and construction monitoring [9].
@@
Earlier researchers implemented decomposition approaches for CPP, including exact cellular [10, 11], trapezoidal [12], Boustrophedon [13], and Morse-based [14] decomposition which divides free space into navigable cells, optimizing systematic path coverage through various cell partitioning and merging techniques. Miao et al. [15] propose a scalable CPP method for cleaning robots using rectangular map decomposition, simplifying the planning process for large-scale applications. These decomposition methods focus on the robot’s physical footprint rather than sensor ranges, such as camera fields of view, which are vital for scanning environments. They create zig-zag patterns and are less suited for patrolling robots, as they might disrupt ongoing activities. Gabriely et al. [16] discuss the Spanning Tree Covering algorithm, one of the first grid-based algorithms for CPP.
@@
Latest methods in coverage path planning (CPP) move beyond merely partitioning areas and integrate both operational aspects and sensor capabilities into the area decomposition process [17, 18, 19, 8, 20, 21, 22, 23, 24]. For instance, Swain et al. [17] introduced an efficient path-planning algorithm for 2D ground area coverage using multi-UAVs, emphasizing the integration of sensor capabilities into the planning process. Mansouri et al. [18] developed a method for 2D visual area coverage and path planning that couples the robot’s path with the camera footprints, enhancing the fidelity of CPP by considering visual sensor constraints. Similarly, Ghaddar et al. [19] presented an energy-aware grid-based CPP for UAVs, incorporating area partitioning in the presence of no-fly zones, which highlights the importance of operational constraints in path planning. Phung et al. [8] proposed an enhanced discrete particle swarm optimization algorithm for UAV vision-based surface inspection, demonstrating the benefits of considering sensor footprints for efficient inspection tasks. Kantaros et al. [23] explored distributed coverage control for concave areas, considering visibility sensing constraints, further highlighting the need to integrate sensor capabilities into the CPP process. An et al. [25] introduced the Rainbow Coverage Path Planning (RCPP) method for patrolling mobile robots with circular sensing ranges, optimizing the path to minimize overlap in sensing regions and enhance patrolling efficiency.
II-B Advancements in Coverage Path Planning approaches
Tran et al. [22, 21, 24] contributed to the dynamic and robust multi-robot coverage of both known and unknown environments through frontier-led swarming techniques. These methods underscore the importance of dynamic adaptation and sensor integration in CPP. Baras and Dasygenis [20] focused on energy-efficient CPP for UGVs by reducing turns, crucial for prolonging operational durations and enhancing efficiency. Jonnarth et al. [26] proposed a deep reinforcement learning approach where the CPP problem is formulated as a Markov Decision Process. Talavera et al. [27] developed an autonomous ground robot to support firefighters in indoor emergencies. This robot utilizes advanced CPP algorithms to navigate complex indoor environments, detect fires, and provide real-time situational awareness to firefighting teams, demonstrating the practical utility of CPP in critical applications. Nasirian et al. [28] addressed the challenge of effectively covering and disinfecting an environment with mobile robots using a graph-based representation. Similarly, Pierson et al. [29] focused on deploying a mobile UVC disinfection robot with an optimized CPP algorithm. Perminov et al. [30] presented a genetic-based human-aware CPP algorithm for autonomous disinfection robots, achieving significant coverage efficiency. Chen et al. [31] utilize building information modeling (BIM) to enhance the efficiency of CPP for indoor robots, considering various robotic configurations for adaptable planning. This work builds on previous research [32] which explored global path planning based on BIM and physics engines for UGVs in indoor environments. Full coverage planning by Lee et al. [33] researches intricacies of path planning in environments with negative obstacles.
II-C Established practical implementations of Coverage Path Planning
@@
Building upon the significant advancements in coverage path planning (CPP) techniques, certain studies have not only concentrated on theoretical models but have also ventured into practical implementations and analyses through software tools. The early work of Bormann et al. [34] an exploration of room segmentation proposed a set of suitable map division algorithms. Following this foundational work, they presented a comprehensive survey alongside the implementation and analysis of various indoor CPP methods [35].
@@ III Proposed Approach: Fast-Revisit Coverage Path Planning (FaRe-CPP)
In FaRe-CPP, instead of considering the robot’s physical dimensions as the footprint, we utilize the field of view (FoV) of the range sensor, e.g. 2D camera, mounted on the robot as the footprint. The range within which the sensor can perceive range information is referred to as the sensor range. The grid cells within the FoV are then referred to as explored cells, and the center cell is considered the robot’s position (x, y) while the orientation () is obtained based on the direction of the robot relative to the coordinate system of the grid map. As visually summarized in Fig. 2, there are three main steps in FaRe-CPP which we describe in the next subsections.
@@ III-B Step 2: Coverage Area Estimation per Waypoint Candidate
where and represent the cell indices in the grid map , represents the side length of each quadratic cell so that indicates the physical area of each cell, and represents the explored area of each sub-grid map with the respective waypoint candidate selected as next actual waypoint.
@@ IV-D Comparison with Existing Algorithms
| Algorithm | Computation Time (s) | Path Length (m) | Total Rotation (rad) | Coverage Area (%) | Revisit Time (s) |
@@
| Proposed FaRe-CPP | 109 | 148.00 1.66 | 227 0 | 95 0 | 948 5.53 |
In this section, we demonstrate our proposed FaRe-CPP’s effectiveness by comparing it to other algorithms that utilize camera footprints qualitatively in Fig. 7 and quantitatively in Table II. As a baseline, we compare our work to the algorithms thoroughly investigated in Bormann et al.’s survey [35] in their well-established, freely available implementations444https://github.com/ipa320/ipa_coverage_planning. We applied these algorithms to our chosen environment grid map, maintaining consistency with their work by using the same parameters for the field of view. Specifically, we considered an RGB-D camera with a horizontal field of view of at the robot’s position and a sensor radius of . The revisit time is determined as explained above with and .
Based on the results shown in Fig. 7, our proposed FaRe-CPP algorithm achieves more efficient coverage with fewer overlaps and a more streamlined path compared to the other methods. Table II quantitatively supports these findings, where we present averages for 10 comparison runs. FaRe-CPP exhibits the shortest path length, lowest total cumulative rotation, and fastest revisit time while maintaining a high coverage area percentage. FaRe-CPP showing no variance in computation time, total rotation and coverage area is based on the fact that these metrics are calculated in the the heuristic, deterministic parts of the algorithm. In total, these metrics highlight the performance of FaRe-CPP in achieving near-optimal coverage with fast revisit times for subsequent patrolling, which is an essential requirement for the targeted use cases.
With respect to computing time, FaRe-CPP currently performs worse than the other algorithms in their respectively used implementation except for Grid TSP. The main reason is that the currently available implementation of our approach uses Python where all others are implemented in C++. While FaRe-CPP works heuristically and often finds near-optimal solutions, its computation time can be exponential, especially as the size and complexity of the grid map increase.
@@ V Conclusion
We evaluated our approach in a simulation environment, demonstrating that our algorithm outperforms existing approaches by achieving optimal coverage areas with shorter path lengths, lower cumulative rotation, and faster revisit times. Compared to the average revisit times and path lengths of existing approaches, FaRe-CPP approximately showed a 45% and 40% reduction, respectively, in these highly relevant performance indicators.

### v2

@@ Abstract
In this paper, we propose the novel Fast-Revisit Coverage Path Planning (FaRe-CPP) algorithm using a greedy heuristic approach to propose waypoints for maximum coverage area and a random search-based path optimization technique to obtain a path along the proposed waypoints with minimum revisit time. We evaluated the algorithm in a simulated environment using Gazebo and a camera-equipped TurtleBot3 against a number of existing algorithms. Compared to their average path lengths and revisit times, our FaRe-CPP algorithm showed a reduction of at least 21% and 33%, respectively, in these highly relevant performance indicators.
@@
Ensuring minimum revisit times of any given waypoint is crucial for certain application, for instance, safety patrol on hazardous substance sites like garbage dumps and recycling centers. Typically, in such sites, spontaneous fires are an imminent hazard because of inflammable substances or self-ignition because of the heat produced during composting. Therefore, fire incidents happen frequently [3, 4] and there is great interest by waste disposal enterprises to prevent such incidents. A safety patrol robot that is able to detect fires autonomously, and potentially even applies extinguishing agents directly, calls the fire department and keeps the fire at bay until their arrival. For such kind of patrols, minimum revisit times with maximum coverage are essential to prevent fires from getting out of control.
In this paper, we present a novel Coverage Path Planning algorithm with focus on fast revisit times for a single mobile patrol robot. We name the algorithm Fast-Revisit Coverage Path Planning (FaRe-CPP). Utilizing a pre-recorded 2D occupancy grid map, our approach updates the surveilled region within the field of view of the relevant onboard sensors at each step, subsequently creating waypoints that maximize the coverage area. Initially generated with a greedy heuristic for maximum perception potential, these waypoints are optimized using the GRASP method [5] to minimize path length while preserving coverage information. Extending existing work from the literature, our approach ensures minimal revisit times as its key contribution. This is achieved by a combination of waypoint generation and optimization which results in maximum overall area coverage with no path overlap and minimal path length.
A Python implementation of FaRe-CPP with ROS111Robot Operating System, https://ros.org connectivity for rapid deployment in real-life scenarios is publicly available on https://github.com/hcr-lab/fare-cpp.
@@ II Related Work
II-A Basic Coverage Path Planning Methods and Applications
@@
Coverage Path Planning (CPP) is essential for robot applications like vacuum cleaning [6], radiation monitoring [7], bridge inspection [8], emergency evacuation [3], and construction monitoring [9].
@@
Earlier researchers implemented decomposition approaches for CPP, including exact cellular [10, 11], trapezoidal [12], Boustrophedon [13], and Morse-based [14] decomposition which divides free space into navigable cells, optimizing systematic path coverage through various cell partitioning and merging techniques. Miao et al. [15] propose a scalable CPP method for cleaning robots using rectangular map decomposition, simplifying the planning process for large-scale applications. These decomposition methods focus on the robot’s physical footprint rather than sensor ranges, such as camera fields of view, which are vital for scanning environments and therefore an essential part of FaRe-CPP. Decomposition methods create zig-zag patterns and are less suited for patrolling robots, as they might disrupt ongoing activities. Gabriely et al. [16] discuss the Spanning Tree Covering algorithm, one of the first grid-based algorithms for CPP.
@@
Latest methods in Coverage Path Planning move beyond merely partitioning areas and integrate both operational aspects and sensor capabilities into the area decomposition process [17, 18, 19, 8, 20, 21, 22, 23, 24]. For instance, Swain et al. [17] introduced an efficient path-planning algorithm for 2D ground area coverage using multi-UAVs, emphasizing the integration of sensor capabilities into the planning process. Mansouri et al. [18] developed a method for 2D visual area coverage and path planning that couples the robot’s path with the camera footprints like our approach, enhancing the fidelity of CPP by considering visual sensor constraints. Similarly, Ghaddar et al. [19] presented an energy-aware grid-based CPP for UAVs, incorporating area partitioning in the presence of no-fly zones, which highlights the importance of operational constraints in path planning. Phung et al. [8] proposed an enhanced discrete particle swarm optimization algorithm for UAV vision-based surface inspection, demonstrating the benefits of considering sensor footprints for efficient inspection tasks. Kantaros et al. [23] explored distributed coverage control for concave areas, considering visibility sensing constraints, further highlighting the need to integrate sensor capabilities into the CPP process. An et al. [25] introduced the Rainbow Coverage Path Planning (RCPP) method for patrolling mobile robots with circular sensing ranges, optimizing the path to minimize overlap in sensing regions and enhance patrolling efficiency.
However, none of the mentioned methods considers fast revisit times for time-critical patrol operations as is the focus of our approach.
II-B Established Practical Implementations of Coverage Path Planning
@@
Building upon the significant advancements in Coverage Path Planning techniques, certain studies have not only concentrated on theoretical models but have also ventured into practical implementations and analyses through software tools. The early work of Bormann et al. [34] provides an exploration of segmented rooms with a set of suitable map division algorithms. Following this foundational work, they presented a comprehensive survey alongside the implementation and analysis of various indoor CPP methods [35]. We are using several of the analyzed methods as baseline for our approach, as described in detail in Section IV-D.
II-C Coverage Path Planning for Environmental Operations
Tran et al. [22, 21, 24] contributed to the dynamic and robust multi-robot coverage of both known and unknown environments through frontier-led swarming techniques. These methods underscore the importance of dynamic adaptation and sensor integration in CPP like our camera-based approach. Talavera et al. [27] developed an autonomous ground robot to support firefighters in indoor emergencies. This robot utilizes advanced CPP algorithms to navigate complex indoor environments, detect fires, and provide real-time situational awareness to firefighting teams, demonstrating the practical utility of CPP in critical firefighting applications like we target with our work. Nasirian et al. [28] addressed the challenge of effectively covering and disinfecting an environment with mobile robots using a graph-based representation like we use in our approach. Similarly, Pierson et al. [29] focused on deploying a mobile UVC disinfection robot with an optimized CPP algorithm. Perminov et al. [30] presented a genetic-based human-aware CPP algorithm for autonomous disinfection robots, achieving significant coverage efficiency. These hygiene-optimizing robot applications are similar to our firefighting approach in the sense that they profit from fast revisit times and are deployed on robots that are to treat the environment.
@@ III Proposed Approach: Fast-Revisit Coverage Path Planning (FaRe-CPP)
In FaRe-CPP, instead of considering the robot’s physical dimensions as the footprint, we utilize the field of view (FoV) of the range sensor, e.g. 2D camera, mounted on the robot as the footprint. The range within which the sensor can perceive range information is referred to as the sensor range. The grid cells within the FoV are then referred to as explored cells, and the center cell is considered the robot’s position (x, y) while the orientation () is obtained based on the direction of the robot relative to the coordinate system of the grid map. As visually summarized in Fig. 2, there are four main steps in FaRe-CPP which we describe in the next subsections.
@@ III-B Step 2: Coverage Area Estimation per Waypoint Candidate
where and represent the cell indices in the sub grid map , represents the side length of each quadratic cell so that indicates the physical area of each cell, and represents the explored area of each sub-grid map with the respective waypoint candidate selected as next actual waypoint.
@@ III-C Step 3: Waypoint Selection by Maximum Coverage Area
Our waypoint selection strategy ensures that at every iteration, we select a waypoint globally from all possible candidates. By evaluating coverage potential over the entire set of waypoint candidates, we prevent the risk of the algorithm converging to a suboptimal local minimum.
@@ IV-D Comparison with Existing Algorithms
(Bormann et al. [35] parameters)
(optimized for long-range efficiency)
Algorithm | Comp. Time (s) | Path Length (m) | Total Rotation (rad) | Coverage Area (%) | Revisit Time (s) |
@@
Proposed FaRe-CPP | 274 5.89 | 199.65 5.06 | 272 13.63 | 92 0 | 1188 28.35 |
(Bormann et al. [35] parameters) | |||||
Proposed FaRe-CPP | 109 3.26 | 148.00 3.63 | 141 5.91 | 95 0 | 764 15.29 |
(optimized for long-range efficiency) |
We evaluate the effectiveness of our proposed FaRe-CPP algorithm through qualitative comparisons (Fig. 7) and quantitative metrics (Table II) against existing camera foot print-based methods. As baselines, we utilize the algorithms evaluated in Bormann et al. [35] as implemented in their publicly available framework444https://github.com/ipa320/ipa_coverage_planning, which represent the state of the art in coverage path planning. For consistency and fairness, we adopt the sensor parameters specified in their work: an RGB-D sensor mounted at 65 cm height, facing downward, with a trapezoidal field of view (FoV) with a 70 cm (closest edge, 15 cm from the robot center) to 1.3 m (farthest edge, 1.15 m from the robot center). This setup yields a horizontal FoV of , a vertical FoV of , and a sensor range of 1.3 m. In addition, we evaluate the optimal parameter set for our approach, which assumes a sensor parallel to the floor, producing a conical FoV constrained by a horizontal field of view and sensing range of 1.3 m . To enable a direct comparison, we evaluate FaRe-CPP under these two configurations: (1) with parameters matching Bormann et al.’s [35] setup, and (2) optimized for long-range efficiency, our primary objective.
As the results show in Fig. 7, FaRe-CPP achieves more efficient coverage in both configurations with fewer overlaps and a more streamlined path compared to the other methods. Table II quantitatively supports these findings, which shows averages for 10 comparison runs. FaRe-CPP exhibits the shortest path length, lowest total cumulative rotation, and fastest revisit time while maintaining a high coverage area percentage.
In total, compared to the average of the baseline approaches, our algorithm achieves a reduction in path length of 21% for similar and 42% for long range-optimized parameter settings. This results in a reduced revisit time of 33% for similar and 57% for long range-optimized parameter settings. These metrics highlight the performance of FaRe-CPP in achieving near-optimal coverage with fast revisit times for subsequent patrolling, which is an essential requirement for the targeted use cases.
With respect to computing time, FaRe-CPP currently lags behind other algorithms except for Grid TSP. To ensure consistent comparison, all methods use the same GRASP-based waypoint optimization, implemented uniformly in Python. The difference lies in waypoint generation, which is implemented in Python for FaRe-CPP and in C++ for the baselines. However, since this stage is heuristic and has minimal effect on final path quality, the impact is limited to computation time, not algorithmic efficiency.
@@ V Conclusion
We evaluated our approach in a simulation environment, demonstrating that our algorithm outperforms existing approaches by achieving optimal coverage areas with shorter path lengths, lower cumulative rotation, and faster revisit times. Compared to the average path lengths and revisit times of existing approaches, FaRe-CPP showed a reduction of at least 21% and 33%, respectively, in these highly relevant performance indicators.